Chris cuts half a rectangle lawn, 40m by 30m, by mowing strips of equal width around the perimeter. Jonah cuts the small rectangle left. How wide a strip does Chris cut so that they share the work equally?
Answer by checkley75(3666)

(40-2X)(30-2X)=(40*30)/2
1200-60X-80X+4X^2=1200/2
1200-140X+4X^2=600
4X^2-140X+1200-600=0
4X^2-140X+600=0
(2X-10)(2X-60)=0
2X-10=0
2X=10
X=5M IS THE WIDTH OF THE STRIP THAT REDUCES TO AREA IN 1/2.
STARTING AREA=40*30=1200
(40-10)(30-10)=1200/2
30*20=600
600=600 1/2 OF ORIGINAL AREA.